OpenStudy (anonymous):

Find the exact value of (Theta) if possible without using a calculator: (Theta) = sin^-1(Tan(60*))

OpenStudy (anonymous):

\[\Theta=\sin^{-1} (\tan (60))\]

jhonyy9 (jhonyy9):

do you know tan(60) = ? so than not you can rewrite it like (sin60)/(cos60) = ? hope you know sin60 and cos60 = ?

OpenStudy (anonymous):

yeah the answer using the calc. is .3257721186

jhonyy9 (jhonyy9):

ok so without calculator sin60 = sqrt3 /2 cos60 = 1/2

jhonyy9 (jhonyy9):

so than tan60 = (sqrt3 /2)/(1/2) = (sqrt3 /2) *2 = sqrt3 yes ?

OpenStudy (anonymous):

yes that it! thank you!!

jhonyy9 (jhonyy9):

so was my pleasure than these will help you good luck bye bye